OpenStudy (anonymous):

Please help. I will give medal. At a bicycle warehouse, the ratio of mountain bikes to street bikes was 3 to 5. If the warehouse had 57 mountain bikes, how many street bikes did it have?

OpenStudy (shamil98):

3/5 = 57/x cross multiply and solve for x.

OpenStudy (anonymous):

So is the answer x? or 95

OpenStudy (jack1):

3 mountain bikes to every 5 street bikes \[\huge \frac {3mtn}{5street} = \frac {57mtn}{x ~ street}\] \[\huge x ~ street= \frac {57mtn}{\frac {3mtn}{5street}}\] \[\huge x ~ street= 57mtn \times \frac {5street}{3mtn}\]

OpenStudy (anonymous):

Sooooooooooo whats the answer? Is the answer 95?

OpenStudy (jack1):

yep, perfect, 95 street

OpenStudy (anonymous):

yay

OpenStudy (anonymous):

use sofer x

OpenStudy (anonymous):